# CogniAlign: Word-Level Multimodal Speech Alignment with Gated Cross-Attention for Alzheimer’s Detection

### v1

@@ Abstract
Early detection of cognitive disorders such as Alzheimer’s disease is critical for enabling timely clinical intervention and improving patient outcomes. In this work, we introduce CogniAlign, a multimodal architecture for Alzheimer’s detection that integrates audio and textual modalities, two non-intrusive sources of information that offer complementary insights into cognitive health. Unlike prior approaches that fuse modalities at a coarse level, CogniAlign leverages a word-level temporal alignment strategy that synchronizes audio embeddings with corresponding textual tokens based on transcription timestamps. This alignment supports the development of token-level fusion techniques, enabling more precise cross-modal interactions. To fully exploit this alignment, we propose a Gated Cross-Attention Fusion mechanism, where audio features attend over textual representations, guided by the superior unimodal performance of the text modality. In addition, we incorporate prosodic cues, specifically interword pauses, by inserting pause tokens into the text and generating audio embeddings for silent intervals, further enriching both streams. We evaluate CogniAlign on the ADReSSo dataset, where it achieves an accuracy of 90.36%, outperforming existing state-of-the-art methods. A detailed ablation study confirms the advantages of our alignment strategy, attention-based fusion, and prosodic modeling.
Index Terms:
Cognitive Decline, Deep Learning, Multimodal, Speech Signal Processing, Natural Language Processing
@@ I Introduction
In response to these needs, we introduce CogniAlign, a multimodal model for Alzheimer’s disease detection that leverages audio and text, two complementary and non-intrusive modalities, to capture indicative patterns of cognitive decline. CogniAlign yields an accuracy of 90.36% on the ADReSSo dataset [23], outperforming existing state-of-the-art methods. Unlike prior work that often treats modalities independently and fuses them only at later stages, our approach performs a fine-grained word-level alignment between audio and text embeddings, allowing for more coherent and semantically grounded fusion. To enable this, we use the Whisper model [24] to transcribe the audio recordings and extract word-level timestamps. These timestamps guide the temporal alignment between textual tokens and audio embeddings.
Additionally, we enrich the model with prosodic information by identifying speech pauses and inserting corresponding pause tokens into the transcription, which are then aligned with audio features to enhance the multimodal representation and boost both unimodal and multimodal performance. To enable effective multimodal fusion, we propose a gated cross-attention fusion strategy combined with a word-level alignment mechanism, ensuring that tokens from each modality correspond to the same speech segments. This approach significantly enhances the model’s ability to capture cross-modal interactions and improves overall performance. The implementation code is publicly available on GitHub  [^1] https://github.com/davidorp/CogniAlign.
@@
The remainder of this paper is organized as follows: Section II reviews the relevant literature on the topic; Section III describes the methodology employed throughout this work; Section IV presents the experimental setup and results; finally, Section V outlines the conclusions drawn from this study.
@@ III-A Textual Encoder
Specifically, textual features are extracted using the frozen DistilBERT model, selected for its performance and computational efficiency, as demonstrated in Section IV-C. The model outputs contextual embeddings for each subword token while preserving the sequence structure from the word-aligned transcription.
@@ III-C Prosodic Information
To capture rhythm and hesitations from natural speech, prosodic information, particularly speech pauses, is extracted and incorporated into the model. These pauses are detected using the word-level timestamps provided by the Whisper model by measuring the temporal gap between consecutive words in the transcription.
@@
In parallel, each detected pause is assigned a corresponding audio embedding. For each pause interval $[s_{p},e_{p}]$, the embedding is computed by averaging the frame-level audio features $f_{j}$ that fall within the pause segment:
@@ III-D Gated Cross-Attention Fusion Strategy
Let $\mathbf{A},\mathbf{T}\in\mathbb{R}^{L\times d}$ denote the audio and textual token sequences, respectively, where $L$ is the number of tokens and $d=768$ is the embedding dimensionality, consistent with the output sizes of DistilBERT and Wav2Vec2. Based on results from our ablation study (Section IV-C), we designate the audio embeddings $\mathbf{A}$ as the queries and the textual embeddings $\mathbf{T}$ as the keys and values in the Transformer cross-attention mechanism:
@@ IV-A Experimental Setup
The experiments in this work are conducted using the ADReSSo dataset [23], specifically designed for Alzheimer’s disease detection. It consists of audio recordings of participants describing the Cookie Theft Picture from the Boston Diagnostic Aphasia Examination [67, 68]. The dataset includes a total of 237 samples, with 118 from individuals diagnosed with Alzheimer’s disease and 119 from cognitively healthy controls. For consistency with the original challenge and related studies [42], we follow a five-fold cross-validation evaluation strategy in our experiments.
The CogniAlign architecture is implemented using a single Transformer Encoder layer with 768 hidden units, 12 attention heads, and an intermediate size of 3072. This configuration aligns with the original architecture of DistilBERT and Wav2Vec2, ensuring compatibility across modalities. All unimodal models, whether textual or audio-based, are equipped with a single Transformer Encoder layer and employ a mean pooling strategy to generate fixed-length representations, maintaining consistency across the architecture.
@@
CogniAlign has been implemented in PyTorch, and all experiments were executed on NVIDIA RTX 4090 GPUs. Models were trained using the AdamW optimizer with a weight decay of 0.1 and an initial learning rate of $2\times 10^{-5}$. A warm-up phase was applied over the first 10 epochs, followed by a cosine annealing learning rate schedule. Training was conducted for a maximum of 200 epochs with a batch size of 32. Early stopping was employed with a patience of 15 epochs to prevent overfitting.
@@ IV-B Comparison with State-Of-The-Art Methods
To assess the effectiveness of the proposed methodology, we compare its performance against current state-of-the-art approaches. Table I reports both accuracy and F1 score results on the ADReSSo dataset, covering audio-based, text-based, and multimodal models. For consistency with prior works, we include results for unimodal settings alongside multimodal approaches, thereby providing a comprehensive view of our model’s relative performance. The unimodal models included correspond to those that achieved the highest performance in our ablation study, as discussed in Section IV-C.
@@
Acc F1 Acc F1 Acc F1
Wang et al., 2021 [73] 75.30 76.00 73.50 73.50 77.20 77.60
Rohanian et al., 2021 [33] 68.00 - 81.00 - 84.00 -
Cui et al., 2023 [30] - 76.30 - 82.60 - 89.40
Ying et al., 2023 [32] 71.20 73.10 78.90 79.00 83.70 83.80
Priyadarshinee et al., 2023 [29] 78.90 - 88.70 - - -
Bang et al., 2024 [31] 69.01 70.39 83.10 83.10 87.32 87.25
Altinok et al., 2024 [42] - - - - 90.00 90.00
CogniAlign (Ours) 80.12 79.46 86.77 86.59 90.36 90.11
TABLE I: Comparison of state-of-the-art models on the ADReSSo dataset across audio, text, and multimodal modalities. Dashes indicate that results were not reported for the respective modality. Bold values denote the best performance.
The results demonstrate the effectiveness of CogniAlign method, surpassing existing state-of-the-art approaches and achieving an accuracy of 90.36% and an F1 score of 90.11% in the multimodal approach. These results are not only promising but also underscore the robustness and reliability of the proposed approach, as it achieves the highest performance across both evaluation metrics. By leveraging aligned and semantically enriched cross-modal representations, the model delivers a more comprehensive understanding of patient communication relevant to Alzheimer’s disease detection.
@@
Regarding the textual modality, our selected model is DistilBERT P, which integrates pause token information into the DistilBERT architecture. This implementation is detailed in subsection IV-C. This model exhibits strong performance, achieving an accuracy of 86.77% and an F1 score of 86.59%. Notably, it obtains the highest F1 score among all compared methods, reflecting a balanced trade-off between precision and recall. Although competitive, this model does not achieve the top accuracy overall. A related study presented in [29], which utilizes the RoBERTa model, reports an accuracy of up to 88.7%. Despite its strong performance in this study, where it was fine-tuned end-to-end, RoBERTa underperformed in our experiments, where it was kept frozen for consistency. This methodological difference likely accounts for the variance in performance outcomes.
@@
For the audio modality, our selected model is Wav2Vec2 P, which utilizes Wav2Vec2 embeddings aligned at the word level and augments them with pause token representations derived from timestamp analysis. The implementation of this approach is detailed in subsection IV-C. This model achieves an accuracy of 80.12% and an F1 score of 79.46%, outperforming all other evaluated audio-based approaches. Its improved performance highlights the benefit of synchronizing audio features with semantic boundaries rather than processing audio as a continuous stream.
@@
IV-C Ablation Study
An extensive ablation study has been performed to identify the most promising and relevant configurations among the proposed approaches. This subsection presents a detailed analysis of the evaluated embeddings, fusion strategies, and key hyperparameters.
@@ IV-C Ablation Study
TABLE IV: Performance of unimodal models with and without prosodic information on five-fold cross-validation using the ADReSSo dataset. Models with names ending in “P” include prosodic features.. Bold values indicate the best overall performance within each modality.
@@
V Conclusions

### v2

@@ Abstract
Early detection of cognitive disorders such as Alzheimer’s disease is critical for enabling timely clinical intervention and improving patient outcomes. In this work, we introduce CogniAlign, a multimodal architecture for Alzheimer’s detection that integrates audio and textual modalities, two non-intrusive sources of information that offer complementary insights into cognitive health. Unlike prior approaches that fuse modalities at a coarse level, CogniAlign leverages a word-level temporal alignment strategy that synchronizes audio embeddings with corresponding textual tokens based on transcription timestamps. This alignment supports the development of token-level fusion techniques, enabling more precise cross-modal interactions. To fully exploit this alignment, we propose a Gated Cross-Attention Fusion mechanism, where audio features attend over textual representations, guided by the superior unimodal performance of the text modality. In addition, we incorporate prosodic cues, specifically interword pauses, by inserting pause tokens into the text and generating audio embeddings for silent intervals, further enriching both streams. We evaluate CogniAlign on the ADReSSo dataset, where it achieves an accuracy of 87.35% over a Leave-One-Subject-Out setup and of 90.36% over a 5 fold Cross-Validation, outperforming existing state-of-the-art methods. A detailed ablation study confirms the advantages of our alignment strategy, attention-based fusion, and prosodic modeling. Finally, we perform a corpus analysis to assess the impact of the proposed prosodic features and apply Integrated Gradients to identify the most influential input segments used by the model in predicting cognitive health outcomes.
@@ I Introduction
In response to these needs, we introduce CogniAlign, a multimodal model for Alzheimer’s disease detection that leverages audio and text, two complementary and non-intrusive modalities, to capture indicative patterns of cognitive decline. CogniAlign yields an accuracy of 87.35% over a Leave-One-Subject-Out (LOSO) and of 90.36% over a 5 fold Cross-Validation (CV) on the ADReSSo dataset [23], outperforming existing state-of-the-art methods. Unlike prior work that often treats modalities independently and fuses them only at later stages, our approach performs a fine-grained word-level alignment between audio and text embeddings, allowing for more coherent and semantically grounded fusion. To enable this, we use the Whisper model [24] to transcribe the audio recordings and extract word-level timestamps. These timestamps guide the temporal alignment between textual tokens and audio embeddings.
Additionally, we enrich the model with prosodic information by identifying speech pauses and inserting corresponding pause tokens into the transcription, which are then aligned with audio features to enhance the multimodal representation and boost both unimodal and multimodal performance. To enable effective multimodal fusion, we propose a gated cross-attention fusion strategy combined with a word-level alignment mechanism, ensuring that tokens from each modality correspond to the same speech segments. This approach significantly enhances the model’s ability to capture cross-modal interactions and improves overall performance. Finally, we conducted a corpus analysis to examine the influence of the proposed prosodic features within the dataset, and we analyzed the Integrated Gradients to identify the most relevant segments of each sample that the model leverages when predicting Alzheimer’s disease and cognitive health. The implementation code is publicly available on GitHub  [^1] https://github.com/davidorp/CogniAlign.
@@
  • We perform a corpus analysis to examine the influence of the proposed prosodic features within the dataset. Additionally, we analyzed the Integrated Gradients to identify the most relevant segments of each sample that the model considers when predicting Alzheimer’s disease and cognitive health.
The remainder of this paper is organized as follows: Section II reviews the relevant literature on the topic; Section III describes the methodology employed throughout this work; Section IV presents the experimental setup and results; Section V introduces an analysis of the corpus and different explainability methods; Section VI discusses the main limitations regarding data availability of this work; finally, Section VII outlines the conclusions drawn from this study.
@@ III-A Textual Encoder
Specifically, textual features are extracted using the frozen DistilBERT model, selected for its performance and computational efficiency, as demonstrated in Section IV-E. The model outputs contextual embeddings for each subword token while preserving the sequence structure from the word-aligned transcription.
@@ III-C Prosodic Information
To capture rhythm and hesitations from natural speech, prosodic information, particularly speech pauses, is extracted and incorporated into the model. These pauses are detected using the word-level timestamps provided by the Whisper model by measuring the temporal gap between consecutive words in the transcription. This process is visually summarized in Figure 2.
Figure 2: Prosodic augmentation pipeline. Pauses, detected using Whisper word-level timestamps (and also visible as silent regions in the waveform), are inserted into the transcription as punctuation marks (comma, period, ellipsis) based on duration. Inserted pauses, shown in green, enhance the original transcription with prosodic cues.
@@
In parallel, each detected pause is assigned a corresponding audio embedding. For each pause interval $s_{p},e_{p}]$, the embedding is computed by averaging the frame-level audio features $f_{j}$ that fall within the pause segment:
@@ III-D Gated Cross-Attention Fusion Strategy
Let $\mathbf{A},\mathbf{T}\in\mathbb{R}^{L\times d}$ denote the audio and textual token sequences, respectively, where $L$ is the number of tokens and $d=768$ is the embedding dimensionality, consistent with the output sizes of DistilBERT and Wav2Vec2. Based on results from our ablation study (Section IV-E), we designate the audio embeddings $\mathbf{A}$ as the queries and the textual embeddings $\mathbf{T}$ as the keys and values in the Transformer cross-attention mechanism:
@@ IV-A Experimental Setup
Dataset: The experiments in this work are conducted using the ADReSSo dataset [23], specifically designed for Alzheimer’s disease detection. It consists of audio recordings of participants describing the Cookie Theft Picture from the Boston Diagnostic Aphasia Examination [67, 68]. The dataset includes a total of 237 samples, with 118 from individuals diagnosed with Alzheimer’s disease and 119 from cognitively healthy controls.
Implementation Details: The CogniAlign architecture is implemented using a single Transformer Encoder layer with 768 hidden units, 12 attention heads, and an intermediate size of 3072. This configuration aligns with the original architecture of DistilBERT and Wav2Vec2, ensuring compatibility across modalities. All unimodal models, whether textual or audio-based, are equipped with a single Transformer Encoder layer and employ a mean pooling strategy to generate fixed-length representations, maintaining consistency across the architecture.
@@
Training Details: CogniAlign has been implemented in PyTorch, and all experiments were executed on NVIDIA RTX 4090 GPUs. Models were trained using the AdamW optimizer with a weight decay of 0.1 and an initial learning rate of $2\times 10^{-5}$. A warmup phase was applied over the first 20 epochs, followed by a cosine learning rate schedule.
Evaluation Details: To identify the most effective combination of modalities and hyperparameters, we conducted an ablation study using 5-fold cross-validation. This training setup ran for a maximum of 200 epochs with a batch size of 32, and early stopping was applied with patience of 15 epochs to mitigate overfitting.
The hyperparameter configuration selected in this ablation stage has been unchanged in every subsequent experiment, covering the regression task and the additional evaluation on the ADReSS dataset.
For comparison with state‑of‑the‑art methods, we report metrics under both 5‑fold Cross‑Validation (CV) and LOSO setups. The LOSO configuration employed the same training parameters as the 5‑CV setup but used a fixed 60‑epoch schedule without early stopping, as no validation set was available. Notably, most prior studies use varying cross‑validation strategies, ranging from 70/30 train–test splits to different k‑fold configurations, which can lead to inconsistencies in model evaluation [32].
To address these discrepancies, we follow the baseline approach [23], adopting LOSO as our primary protocol for comparison because it guarantees replicability and ensures a more robust, fair evaluation. Beyond mitigating validation inconsistencies, LOSO performs inference on each subject independently, ensuring that every test sample is entirely isolated from the training data. Unlike traditional k‑fold cross‑validation, LOSO avoids arbitrary data splits and eliminates subject overlap between training and testing sets, providing a more realistic assessment of a model’s generalisability to previously unseen individuals. Meanwhile, the 5‑CV protocol is retained for the ablation study owing to its lighter computational cost, which allows rapid, exhaustive exploration of modality combinations and hyper‑parameter settings.
@@ IV-B Comparison with State-Of-The-Art Methods
To assess the effectiveness of the proposed methodology, we compare its performance against current state-of-the-art approaches. Table I reports accuracy, F1-Score, precision and recall results on the ADReSSo dataset, covering audio-based, text-based, and multimodal models. For consistency with prior works, we include results for unimodal settings alongside multimodal approaches, thereby providing a comprehensive view of our model’s relative performance. The unimodal models included correspond to those that achieved the highest performance in our ablation study, as discussed in Section IV-E.
@@
Protocol Acc F1 Pr Rec Acc F1 Pr Rec Acc F1 Pr Rec
Wang et al., 2021 [73] 10CV 75.30 76.00 77.40 74.70 73.50 73.50 75.30 73.60 77.20 77.60 78.70 74.00
Rohanian et al., 2021 [33] - 68.00 - - - 81.00 - - - 84.00 - - -
Cui et al., 2023 [30] 80/20 - 76.30 - - - 82.60 - - - 89.40 - -
Ying et al., 2023 [32] 10CV 71.20 73.10 73.10 71.00 78.90 79.00 83.10 79.00 83.70 83.60 85.30 84.00
Priyad. et al., 2023 [29] 70/30 78.90 - - - 88.70 - - - - - - -
CogniAlign (Ours) 5CV 80.12 79.46 79.88 81.40 86.77 86.59 87.07 86.83 90.36 90.11 90.15 90.76
Baseline, 2021 [23] LOSO 78.92 - - - 80.72 - - - - - - -
Bang et al., 2024 [31] LOSO 69.01 70.39 70.60 70.42 83.10 83.10 83.10 83.10 87.32 87.25 88.06 87.32
CogniAlign (Ours) LOSO 74.70 76.40 74.73 78.16 82.53 82.63 86.25 79.31 87.35 87.57 90.24 85.06
TABLE I: Comparison of state-of-the-art models on the ADReSSo dataset across audio, text, and multimodal modalities. The top half presents results under various evaluation protocols, including our 5CV setting. The bottom half compares models using the main evaluation metric (LOSO), which enables a fair comparison. Bold values indicate the best performance under LOSO. Dashes indicate unreported evaluation protocols or results. For simplicity, “Acc” denotes accuracy, “Pr” denotes precision, and “Rec” denotes recall.
The results demonstrate the effectiveness of the CogniAlign method, which outperforms existing state-of-the-art approaches. Under the LOSO setup, it achieves an accuracy of 87.35% and an F1 score of 87.57% using the multimodal approach. In the 5-fold CV setup, it attains an accuracy of 90.36% and an F1 score of 90.11%. In this setup, where ablation studies were conducted to optimize unimodal configurations, our CogniAlign unimodal models outperform almost all existing state-of-the-art approaches. More importantly, the multimodal fusion strategy achieves the best performance across both evaluation setups (LOSO and 5-fold CV), even without relying on the strongest unimodal results in the LOSO setup. Regarding the main comparison setup, LOSO, CogniAlign achieves the best multimodal results in all metrics but recall. We attribute the slight drop in recall to the model’s conservative prediction strategy, which prioritizes precision over capturing all positive cases. This highlights the effectiveness of our gated cross-attention mechanism with word-level fusion and underscores the value of integrating multiple modalities in a coordinated manner, rather than treating them as independent sources of information.
In contrast, several existing works adopt less integrated fusion strategies. For instance, the approach presented in [31], as well as those in [32] and [30], rely on simple concatenation of features processed by separate models, each handling input modalities independently. A similar technique is employed in [33], albeit incorporating a gating mechanism to regulate the information flow. Meanwhile, the study in [73] explores attention-based methods. However, it still treats input data from each modality in isolation. Furthermore, the baseline [23] and the method proposed in [29] do not incorporate any form of multimodal fusion. These comparisons underscore the limitations of approaches that fail to fully integrate multimodal information and further validate the strength of our proposed fusion strategy in capturing cross-modal relationships effectively.
@@
Regarding the textual modality, our selected model is DistilBERT P, which integrates pause token information into the DistilBERT architecture. This implementation is detailed in subsection IV-E. This model exhibits strong performance, achieving an accuracy of 82.53% and an F1 score of 82.63% under the LOSO setup, and an accuracy of 86.77% with an F1 score of 86.59% in the 5 CV. Although competitive, this model does not achieve the top accuracy overall. A related study presented in [29], which utilizes the RoBERTa model, reports an accuracy of up to 88.7%. Despite its strong performance in this study, where it was fine-tuned end-to-end, RoBERTa underperformed in our experiments, where it was kept frozen for consistency. This methodological difference likely accounts for the variance in performance outcomes.
@@
For the audio modality, our selected model is Wav2Vec2 P, which utilizes Wav2Vec2 embeddings aligned at the word level and augments them with pause token representations derived from timestamp analysis. The implementation of this approach is detailed in subsection IV-E. This model achieves an accuracy of 76.40% and an F1 score of 76.40% under the LOSO setup, achieving the state-of-the-art performance on F1, precission and recall. In the 5 fold CV, this model achieves an accuracy of 80.12% with an F1 score of 79.46%, outperforming the rest of the proposed methodologies. Its performance highlights the benefit of synchronizing audio features with semantic boundaries rather than processing audio as a continuous stream.
@@
IV-C Regression Task
In addition to a binary label, indicating either cognitive health or Alzheimer’s disease, the ADReSSo dataset also provides a numerical score ranging from 0 to 30. This score corresponds to the result of the Mini-Mental State Examination (MMSE), a widely used clinical tool in the healthcare field for assessing cognitive impairment [74]. The MMSE consists of a series of eleven questions designed to evaluate various cognitive functions. It is important to note that cognitive decline is not a strictly binary condition; rather, it is a progressive process that is more accurately represented on a continuous scale.
To further evaluate our method in estimating the MMSE score, we conducted a regression task. The results, obtained through 5-fold cross-validation, yielded a RMSE of 4.77. Among the works referenced for comparison in the classification task, only the study in [33] reports results for the regression task, achieving an RMSE of 4.26. However, the cross-validation protocol used in that study is not explicitly stated, limiting direct comparability. Additionally, the baseline model proposed in ADReSSo dataset was evaluated using LOSO cross-validation, resulting in a best RMSE of 5.52. Under the same LOSO evaluation protocol, our method achieves an RMSE of 5.28, thereby outperforming the baseline approach.
IV-D Test on external dataset
In addition to the main evaluation presented over the ADReSSo dataset, we further examined the generalizability of our approach using data derived from the DementiaBank corpora. While the Pitt Corpus [27] remains the most widely used benchmark in speech-based dementia research, several of its original recordings contain residual tape noise and other artefacts that may distort acoustic features. To address these limitations, we employed the Alzheimer’s Dementia Recognition through Spontaneous Speech (ADReSS) dataset [75], a carefully curated, acoustically pre-processed, and demographically balanced subset of the Pitt Corpus, and conducted a 5-fold cross-validation experiment. Although ADReSS includes manual transcripts, we regenerated them using the Whisper ASR model to obtain the word-level timestamps required by our alignment module.
CogniAlign achieved an accuracy of 85.06% on this dataset across five-fold cross-validation, demonstrating the method’s generalizability. The current state-of-the-art performance on this dataset is reported in [76], where the authors achieved 90% accuracy. Notably, their approach used the original manual transcriptions, which typically led to improved results, as also shown by the ADReSSo baseline. As with the regression task, no ablation studies or hyperparameter tuning were performed for this evaluation.
IV-E Ablation Study
An extensive ablation study has been performed to identify the most promising and relevant configurations among the proposed approaches. This subsection presents a detailed analysis of the evaluated embeddings, fusion strategies, and key hyperparameters. All experiments presented in this section were conducted using a 5-fold cross-validation setup.
@@ IV-E Ablation Study
TABLE IV: Performance of unimodal models with and without prosodic information on five-fold cross-validation using the ADReSSo dataset. Models with names ending in “P” include prosodic features. Bold values indicate the best overall performance within each modality.
@@
The addition of the gating mechanism not only enhances the performance of the cross-attention model but also yields improvements in its bidirectional variant. To ensure that the observed performance gains from the gated fusion are not attributable to sampling noise, we conducted a paired t-test [92] on 50 accuracy pairs derived from five random seeds combined with ten-fold stratified cross-validation.
The test revealed a mean accuracy The test revealed a mean accuracy improvement of 1.36 percentage points for the gated cross-attention model over the standard cross-attention version, which was statistically significant (p ¡ 0.05). This metric allows us to conclude that the gating mechanism provides a statistically significant advantage over the standard cross-attention approach [93].
@@
V Explainability
In healthcare-related tasks, it is essential to understand how models make decisions, particularly which features or patterns they rely on when predicting Alzheimer’s disease or assessing cognitive health, to ensure explainability [95, 96]. This information can also be valuable for healthcare professionals, offering insights into the model’s reasoning and potentially supporting clinical interpretation. Accordingly, in this section, we investigate the explainability of our approach by conducting a statistical corpus analysis of the ADReSSo dataset and examining the attention mechanisms within the models.
V-A Prosodic information
In this work, we incorporate prosodic cues into the original textual transcriptions by adding pauses. As previously noted, one of the common symptoms of Alzheimer’s disease is word-finding difficulty. This often manifests as hesitations in speech, where the speaker pauses to search for the appropriate word. These pauses are typically represented in transcriptions using commas, periods, or ellipses, depending on their duration [97]. To better capture this behavior, we categorized pauses in the patients’ speech into three types based on their duration. The inclusion of these pause types has led to improved model performance, as demonstrated in the ablation study and supported by findings in related research [63, 64, 33]. Additionally, a corpus-level analysis of these pauses was conducted, with results presented in Table IX.
Pause Original Additional
CH AD CH AD
“,” 5.49 5.17 2.27 1.85
“.” 11.85 13.94 2.67 2.85
“…” 0.29 1.06 3.15 5.85
TABLE IX: Average (mean) number of occurrences per transcription for different pause types. Values are reported separately for cognitively healthy (CH) and Alzheimer’s disease (AD) participants, in both the original transcriptions and the additional prosodic annotations. Bold values indicate the higher mean within each row.
The results reveal a tendency for cognitively healthy individuals to produce shorter pauses, as observed in both the original and the additional transcriptions. These shorter pauses, typically represented by commas, are embedded within sentences and do not necessarily indicate hesitation, as they reflect only brief interruptions in speech.
In contrast, longer pauses are more frequently associated with individuals diagnosed with Alzheimer’s disease. This trend becomes increasingly evident with longer pauses, particularly those marked by ellipses. A similar pattern was reported in [98], where healthy participants exhibited a higher frequency of short pauses. On the other hand, individuals with Alzheimer’s showed a greater tendency toward longer pauses.
These findings are further supported by differences in the average length of audio recordings. In this dataset, the mean duration of audio samples for participants with Alzheimer’s disease is 87.61 seconds, compared to 68.76 seconds for cognitively healthy individuals. This contrast is even more pronounced when considering the number of words per transcription: healthy individuals average 97.16 words per sample, while those with Alzheimer’s average 118.96 words. This suggests that despite producing longer recordings, individuals with Alzheimer’s may require more time and pauses to produce fewer words.
V-B Model Explainability
To better understand the contribution of each input token to the model’s final prediction, we applied Integrated Gradients (IG) [99], a gradient-based attribution method. This approach quantifies the influence of individual input features by integrating the gradients of the model’s output along a linear path from a baseline input (typically a zero vector) to the actual input. In our setup, IG has been employed to compute token-level attribution scores for each prediction, representing relative importance. The results of this analysis are presented in Figure 4.
(a) Influence for the prediction of Alzheimer’s disease.
(b) Influence for the prediction of Alzheimer’s disease.
(c) Influence for the prediction of cognitive healthy.
(d) Influence for the prediction of Alzheimer’s disease.
Figure 4: Integrated Gradients attribution visualization from CogniAlign Model. Tokens contributing positively to the predicted class are shown in green, while those reducing its likelihood are in red. Color intensity corresponds to the relative importance of each token.
Based on the results obtained from the evaluation of the full CogniAlign architecture with prosodic enhancement, we observe a strong influence of prosodic cues on the model’s predictions. Specifically, these cues increase the model’s propensity to predict Alzheimer’s disease while simultaneously reducing its propensity to predict cognitive health. Such markers are commonly associated with individuals experiencing word-finding difficulties or requiring additional time to organize their thoughts, thereby offering both interpretability and clinical relevance.
Regarding lexical features, certain discourse markers such as “okay” and “well” also exhibit a positive influence on the prediction of Alzheimer’s disease. This observation is consistent with our earlier analysis of prosodic cue distributions in the dataset. Similar patterns have been reported in prior studies in the field [100, 101], further validating our findings. These insights not only support the model’s behavior but also align with established clinical observations [102], which report that individuals with Alzheimer’s disease frequently exhibit longer pauses, hesitations, and disrupted discourse structure due to lexical retrieval difficulties. The model’s reliance on prosodic cues such as extended silences and discourse markers like “okay” or “well” reflects these clinical symptoms, thereby enhancing both its interpretability and practical utility in a healthcare context.
VI Limitations
While our methodology, along with other state-of-the-art approaches, demonstrates significant and relevant results, a clear limitation remains in this research field: the lack of large, diverse datasets suitable for this task. The primary reason for this limitation is the sensitivity of the data, particularly audio recordings of individuals with cognitive impairments, which significantly restricts access and sharing.
Creating such datasets is a complex and time-consuming process. In this domain, the challenges are even greater, as data collection must involve healthcare professionals to design and oversee the recording protocols. Moreover, navigating the bureaucratic procedures and obtaining the necessary permissions to handle and share sensitive data, even for academic purposes, presents an additional barrier.
This limitation in dataset size and diversity affects not only the performance of current studies but also the feasibility of deploying these models in real-world scenarios. Small and homogeneous datasets increase the risk of overfitting, where models perform well on evaluation metrics but fail to generalize to broader, real-world populations. Furthermore, many of the existing datasets are collected from narrow demographic groups, often limited to a single language or culture. Without sufficient representation of diverse cultures and languages, the deployment of these algorithms beyond the original context becomes unfeasible.
Therefore, there is a critical need for the development of large-scale, multimodal, and multicultural datasets. Such resources are essential to ensure robust, fair, and generalizable training of deep learning models in this sensitive and impactful area.
VII Conclusions